# MoLA: Motion Generation and Editing with Latent Diffusion Enhanced by Adversarial Training

### v3

@@ Appendix D Evaluation metrics details
We provide more details of evaluation metrics in Section 4.1. We use five metrics to quantitatively evaluate text-to-motion models. These metrics are calculated based on motion and text features extracted with pre-trained networks. More specifically, we utilize the motion encoder and text encoder provided in (Guo et al., 2022). We denote ground-truth motion features, generated motion features, and text features as $f_{\text{gt}}=\mathcal{E}_{\text{M}}(x_{\text{gt}})\in\mathbb{R}^{512}$, $f_{\text{pred}}=\mathcal{E}_{\text{M}}(x_{\text{pred}})\in\mathbb{R}^{512}$, and $f_{\text{text}}=\mathcal{E}_{\text{T}}(c)\in\mathbb{R}^{512}$, where $\mathcal{E}_{\text{M}}(\cdot)$ and $\mathcal{E}_{\text{T}}(\cdot)$ represent the motion encoder and the text encoder, respectively. We explain the five metrics below.
@@
(15) $\text{FID}:=\|\mu_{\text{gt}}-\mu_{\text{pred}}\|^{2}_{2}-\mathrm{tr}(\Sigma_{\text{gt}}+\Sigma_{\text{pred}}-(2\Sigma_{\text{gt}}\Sigma_{\text{pred}})^{\frac{1}{2}}),$
@@
(16) $\text{MMDist}:=\frac{1}{N}\sum_{i=1}^{N}\|f_{\text{pred},i}-f_{\text{text},i}\|_{2}^{2},$
@@
(17) $\text{Diversity}:=\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}\|v_{i}-v^{\prime}_{i}\|_{2}^{2}.$
@@
(18) $\text{MModality}=\frac{1}{K\cdot S_{l}}\sum_{k=1}^{K}\sum_{i=1}^{S_{l}}\|v_{k,i}-v^{\prime}_{k,i}\|_{2}^{2}.$
@@ Appendix E Additional experiments
E.1. Performance comparison on the KIT-ML dataset
We evaluate our method not only on the HumanML3D dataset presented in Section 4.1 but also on the KIT-ML dataset (Plappert et al., 2016), comparing its performance against other text-to-motion generation methods. The KIT-ML dataset includes 3,911 human motion sequences and 6,278 text descriptions derived from the KIT (Mandery et al., 2015) and CMU (CMU, 2003) datasets. These data are split into training, validation, and test sets with proportions of 80%, 5%, and 15%, respectively. Consistent with Section 4.1, the evaluated methods are categorized into three groups: (i) those using VQ-based latent representations (discrete), (ii) those employing data-space diffusion models (continuous, raw data), and (iii) those utilizing VAE-based latent representations (continuous, latent). As shown in Table 4, MoLA achieves the best performance in terms of R-Precision and MMDist, particularly among continuous-based methods. However, it does not outperform methods like MLD in terms of FID. Comparing this with Table 1, we observe a discrepancy in the FID trends across the two datasets, suggesting there is room for improvement in this aspect of MoLA.
@@ E.1. Performance comparison on the KIT-ML dataset
Discrete M2DM (Kong et al., 2023) $0.416^{\pm.004}$ $0.628^{\pm.004}$ $0.743^{\pm.004}$ $0.515^{\pm.029}$ $3.015^{\pm.017}$ $11.417^{\pm.97}$ $3.325^{\pm.37}$
AttT2M (Zhong et al., 2023) $0.413^{\pm.006}$ $0.632^{\pm.006}$ $0.751^{\pm.006}$ $0.870^{\pm.039}$ $3.039^{\pm.021}$ $10.96^{\pm.123}$ $2.281^{\pm.047}$
T2M-GPT (Zhang et al., 2023b) $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.029}$ $10.921^{\pm.108}$ $1.570^{\pm.039}$
MoMask (Guo et al., 2024) $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.131^{\pm.043}$
DiverseMotion (Lou et al., 2023) $0.416^{\pm.005}$ $0.637^{\pm.008}$ $0.760^{\pm.011}$ $0.468^{\pm.098}$ $2.892^{\pm.041}$ $10.873^{\pm.101}$ $2.062^{\pm.079}$
MMM (Pinyoanuntapong et al., 2024b) $0.381^{\pm.005}$ $0.590^{\pm.006}$ $0.718^{\pm.005}$ $0.429^{\pm.019}$ $3.146^{\pm.019}$ $10.633^{\pm.097}$ $1.105^{\pm.026}$
ParCO (Zou et al., 2024) $0.430^{\pm.004}$ $0.649^{\pm.007}$ $0.772^{\pm.008}$ $0.453^{\pm.027}$ $2.820^{\pm.028}$ $10.95^{\pm.094}$ $1.245^{\pm.022}$
BAMM (Pinyoanuntapong et al., 2024a) $0.438^{\pm.009}$ $0.661^{\pm.009}$ $0.788^{\pm.005}$ $0.183^{\pm.013}$ $2.723^{\pm.026}$ $11.008^{\pm.094}$ $1.609^{\pm.065}$
Continuous (raw data) MotionDiffuse (Zhang et al., 2024) $0.417^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $2.958^{\pm.005}$ $11.10^{\pm.143}$ $0.730^{\pm.013}$
MDM (Tevet et al., 2023) $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $0.497^{\pm.021}$ $9.191^{\pm.022}$ $10.847^{\pm.109}$ $1.907^{\pm.214}$
Fg-T2M  (Wang et al., 2023) $0.418^{\pm.005}$ $0.626^{\pm.004}$ $0.745^{\pm.004}$ $0.571^{\pm.047}$ $3.114^{\pm.015}$ $10.93^{\pm.083}$ $1.019^{\pm.029}$
Continuous (latent) MLD (Chen et al., 2023) $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $\bm{0.404^{\pm.027}}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $\bm{2.192^{\pm.071}}$
MotionLCM (Dai et al., 2024) -- -- -- -- -- -- --
@@
Table 4. Comparison with state-of-the-art methods on KIT-ML. Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
E.2. Ablation of classifier-free diffusion guidance on stage 2
As explained in Section 3.3, we employ a classifier-free diffusion guidance technique (Ho and Salimans, 2022). In training, we randomly drop the condition $\tau(\bm{c})$ with a probability of 10% and train both the conditional model $\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))$ and the unconditional model $\epsilon_{\theta}(z_{t},t,\varnothing)$. In inference, the two predictions are linearly combined as follows:
(19) $\epsilon_{\theta}(z_{t},t,\bm{c})=s\cdot\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))+(1-s)\cdot\epsilon_{\theta}(z_{t},t,\varnothing),$
@@ E.2. Ablation of classifier-free diffusion guidance on stage 2
Figure 8. Evaluation sweep over guidance scale $s$

### v4

@@ Appendix D Evaluation metrics details
We provide more details of evaluation metrics in Section 4.1. We use five metrics to quantitatively evaluate text-to-motion models. These metrics are calculated based on motion and text features extracted with pre-trained networks. More specifically, we utilize the motion encoder and text encoder provided in [12]. We denote ground-truth motion features, generated motion features, and text features as $f_{\text{gt}}=\mathcal{E}_{\text{M}}(x_{\text{gt}})\in\mathbb{R}^{512}$, $f_{\text{pred}}=\mathcal{E}_{\text{M}}(x_{\text{pred}})\in\mathbb{R}^{512}$, and $f_{\text{text}}=\mathcal{E}_{\text{T}}(c)\in\mathbb{R}^{512}$, where $\mathcal{E}_{\text{M}}(\cdot)$ and $\mathcal{E}_{\text{T}}(\cdot)$ represent the motion encoder and the text encoder, respectively. We explain the five metrics below.
@@
$\text{FID}:=\|\mu_{\text{gt}}-\mu_{\text{pred}}\|^{2}_{2}-\mathrm{tr}(\Sigma_{\text{gt}}+\Sigma_{\text{pred}}-(2\Sigma_{\text{gt}}\Sigma_{\text{pred}})^{\frac{1}{2}}),$ (13)
@@
$\text{MMDist}:=\frac{1}{N}\sum_{i=1}^{N}\|f_{\text{pred},i}-f_{\text{text},i}\|_{2}^{2},$ (14)
@@
$\text{Diversity}:=\frac{1}{S_{d}}\sum_{i=1}^{S_{d}}\|v_{i}-v^{\prime}_{i}\|_{2}^{2}.$ (15)
@@
$\text{MModality}=\frac{1}{K\cdot S_{l}}\sum_{k=1}^{K}\sum_{i=1}^{S_{l}}\|v_{k,i}-v^{\prime}_{k,i}\|_{2}^{2}.$ (16)
@@ Appendix E Additional experiments
E.1 Performance comparison on the KIT-ML dataset
We evaluate our method not only on the HumanML3D dataset presented in Section 4.1 but also on the KIT-ML dataset [30], comparing its performance against other text-to-motion generation methods. The KIT-ML dataset includes 3,911 human motion sequences and 6,278 text descriptions derived from the KIT [27] and CMU [5] datasets. These data are split into training, validation, and test sets with proportions of 80%, 5%, and 15%, respectively. Consistent with Section 4.1, the evaluated methods are categorized into three groups: (i) those using VQ-based latent representations (discrete), (ii) those employing data-space diffusion models (continuous, raw data), and (iii) those utilizing VAE-based latent representations (continuous, latent). As shown in Table 5, MoLA achieves the best performance in terms of R-Precision and MMDist, particularly among continuous-based methods. However, it does not outperform methods like MLD in terms of FID. Comparing this with Table 1, we observe a discrepancy in the FID trends across the two datasets, suggesting there is room for improvement in this aspect of MoLA.
@@ E.1 Performance comparison on the KIT-ML dataset
Discrete M2DM [21] $0.416^{\pm.004}$ $0.628^{\pm.004}$ $0.743^{\pm.004}$ $0.515^{\pm.029}$ $3.015^{\pm.017}$ $11.417^{\pm.97}$ $3.325^{\pm.37}$
AttT2M [48] $0.413^{\pm.006}$ $0.632^{\pm.006}$ $0.751^{\pm.006}$ $0.870^{\pm.039}$ $3.039^{\pm.021}$ $10.96^{\pm.123}$ $2.281^{\pm.047}$
T2M-GPT [44] $0.416^{\pm.006}$ $0.627^{\pm.006}$ $0.745^{\pm.006}$ $0.514^{\pm.029}$ $3.007^{\pm.029}$ $10.921^{\pm.108}$ $1.570^{\pm.039}$
MoMask [13] $0.433^{\pm.007}$ $0.656^{\pm.005}$ $0.781^{\pm.005}$ $0.204^{\pm.011}$ $2.779^{\pm.022}$ $1.131^{\pm.043}$
DiverseMotion [24] $0.416^{\pm.005}$ $0.637^{\pm.008}$ $0.760^{\pm.011}$ $0.468^{\pm.098}$ $2.892^{\pm.041}$ $10.873^{\pm.101}$ $2.062^{\pm.079}$
MMM [29] $0.381^{\pm.005}$ $0.590^{\pm.006}$ $0.718^{\pm.005}$ $0.429^{\pm.019}$ $3.146^{\pm.019}$ $10.633^{\pm.097}$ $1.105^{\pm.026}$
ParCO [49] $0.430^{\pm.004}$ $0.649^{\pm.007}$ $0.772^{\pm.008}$ $0.453^{\pm.027}$ $2.820^{\pm.028}$ $10.95^{\pm.094}$ $1.245^{\pm.022}$
BAMM [28] $0.438^{\pm.009}$ $0.661^{\pm.009}$ $0.788^{\pm.005}$ $0.183^{\pm.013}$ $2.723^{\pm.026}$ $11.008^{\pm.094}$ $1.609^{\pm.065}$
Continuous (raw data) MotionDiffuse [46] $0.417^{\pm.004}$ $0.621^{\pm.004}$ $0.739^{\pm.004}$ $1.954^{\pm.062}$ $2.958^{\pm.005}$ $11.10^{\pm.143}$ $0.730^{\pm.013}$
MDM [37] $0.164^{\pm.004}$ $0.291^{\pm.004}$ $0.396^{\pm.004}$ $0.497^{\pm.021}$ $9.191^{\pm.022}$ $10.847^{\pm.109}$ $1.907^{\pm.214}$
Fg-T2M  [38] $0.418^{\pm.005}$ $0.626^{\pm.004}$ $0.745^{\pm.004}$ $0.571^{\pm.047}$ $3.114^{\pm.015}$ $10.93^{\pm.083}$ $1.019^{\pm.029}$
Continuous (latent) MLD [3] $0.390^{\pm.008}$ $0.609^{\pm.008}$ $0.734^{\pm.007}$ $\bm{0.404^{\pm.027}}$ $3.204^{\pm.027}$ $10.80^{\pm.117}$ $\bm{2.192^{\pm.071}}$
MotionLCM [6] -- -- -- -- -- -- --
@@
Table 5: Comparison with state-of-the-art methods on KIT-ML dataset. Note that discrete representations do not allow for training-free motion editing; therefore, methods based on VQ-based latent representations (Discrete) are grayed out. The best scores for each metric in the methods using VAE-based latent representations (Continuous (latent)) are highlighted in bold.
E.2 Ablation of classifier-free diffusion guidance on stage 2
As explained in Section 3.3, we employ a classifier-free diffusion guidance technique [15]. In training, we randomly drop the condition $\tau(\bm{c})$ with a probability of 10% and train both the conditional model $\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))$ and the unconditional model $\epsilon_{\theta}(z_{t},t,\varnothing)$. In inference, the two predictions are linearly combined as follows:
$\epsilon_{\theta}(z_{t},t,\bm{c})=s\cdot\epsilon_{\theta}(z_{t},t,\tau(\bm{c}))+(1-s)\cdot\epsilon_{\theta}(z_{t},t,\varnothing),$ (17)
@@ E.2 Ablation of classifier-free diffusion guidance on stage 2
Figure 6: Evaluation sweep over guidance scale $s$